OpenStudy (anonymous):

hello guys i confuse how to clear this question. can someone help me? 1) integral sec^4 (2x) dx 2) sin^3 (3x) cos^5(3x) dx

OpenStudy (anonymous):

\[\int\limits \sec ^4(2x) dx=\int\limits \sec ^2(2x) \sec ^2(2x) dx=\int\limits \left( 1+\tan ^2(2x) \right)\sec ^2(2x)dx\] \[=\int\limits \sec ^2(2x)dx+\int\limits \tan ^2(2x)\sec ^2(2x)dx+c\] \[\int\limits f^n \left( x \right)f \prime \left( x \right)dx=\frac{ f ^{n+1}\left( x \right) }{ n+1 }\] in the second put tan 2x=t,diff.

OpenStudy (anonymous):

2.\[\int\limits \sin ^3 x \cos ^5x~dx=\int\limits \cos ^5x \sin ^2x \sin x~dx\] \[=\int\limits \cos ^5x~\left( 1-\cos ^2x \right)\sin x~dx=\int\limits \cos ^5x ~\sin x~dx-\int\limits \cos ^7x~\sin x~dx+c\] solve it.